# On $\mathcal{F}$-semi-transitive families of operators

### v2

@@ Abstract.
Motivated by the concept of disjoint topological transitivity for supercyclicity given in [30], in this paper we introduce the concept of disjoint Furstenberg-semi-transitivity. Then we provide sufficient conditions for cosine operator functions generated by adjoints of weighted composition operators to be Furstenberg-semi-transitive on the space of Radon measures on a locally compact, non-compact Hausdorff space. Also, we characterize disjoint Furstenberg-semi-transitive generalized bilateral weighted shifts on the standard Hilbert module over the C*-algebra of compact operators on a separable Hilbert space. As a special case, we consider the case when the corresponding Furstenberg family under consideration is the family of all infinite subsets of the set of natural numbers. Finally, we illustrate our results with concrete examples.
@@ 1. Introduction
The research on the dynamics of weighted composition operators has started from pioneering papers by Salas on the dynamics on weighted shifts, see [33, 35]. Since then, numerous papers on the dynamics of weighted composition operators on various function spaces have been published. In particular, the dynamics of cosine operator functions generated by weighted composition operators has been studied in for instance [10, 11, 12, 13, 14, 23, 37]. Moreover, the dynamics of cosine operator functions generated by some other operators has been investigated in for instance [22, 27, 28]. Further, Furstenberg topological transitivity of weighted composition operators, as well as of some other classes of operators, has been studied in several papers, see [2, 7, 18, 19, 29]. Finally, quite recently, topological transitivity of the adjoints of weighted composition operators has been investigated and studied in [22]. These adjoints considered in [22] act on the space of Radon measures on a locally compact, non-compact Hausdorff space. All these facts motivated us to study in this paper the Furstenberg-dynamics of cosine operator functions generated by the adjoints of weighted composition operators. In Section 3 in this paper, we introduce first the concept of $\mathcal{F}$-semi-transitivity, as a generalization of the concept of topological transitivity for supercyclicity given in [30]. Then, in Theorem 3.1, we obtain sufficient conditions for cosine operator functions generated by the adjoints of weighted composition operators to be $\mathcal{F}$-semi-transitive on the space of Radon measures (on a locally compact, non-compact Hausdorff space) equipped with the total variation norm. A special case of this theory is the case when the corresponding Furstenberg family under consideration is the family of all infinite subsets of $\mathbb{N},$ as given in Corollary 3.2. We provide also concrete examples where the sufficient conditions of Theorem 3.1 and Corollary 3.2 in this paper are satisfied, but where the sufficient conditions of [22, Proposition 3.3] are not satisfied, illustrating thus the difference between (ordinary) topological transitivity and $\mathcal{F}$-semi-transitivity of cosine operator functions.
@@ 2. Preliminaries
    If $A$ belongs to $\mathcal{F}$ and $B$ contains $A$, then $B$ also belongs to $\mathcal{F}$.
@@ 3. $\mathcal{F}$-semi-transitive cosine operator functions
Assume that $\mathcal{S}$ is a set and $\mathcal{F}$ is a proper Furstenberg family of subsets of $\mathcal{S}$. Let $(\alpha_{z})_{z\in\in\mathcal{S}}$ be a family of homeomorphisms of $\Omega$ and $(w_{z})_{z\in\in\mathcal{S}}$ be a family of positive bounded continuous functions on $\Omega$ such that $\frac{1}{w_{z}}$ is also bounded for all $z\in\mathcal{S}$. For each $z\in\mathcal{S}$, we let $T_{z}^{\ast}:=T_{\alpha_{z},w_{z}}^{\ast}$ and let $C_{z}^{\ast}:=\frac{1}{2}(T_{z}^{\ast}+S_{z}^{\ast})$. We will say that $\{C_{z}^{\ast}\}_{z\in\mathcal{S}}$ is $\mathcal{F}$-semi-transitive if for every non-empty open subsets $U_{1}$ and $U_{2}$ of $M(\Omega)$, the set

### v3

@@ Abstract.
Motivated by the concept of disjoint topological transitivity for supercyclicity given in [30], in this paper, we introduce the concept of disjoint Furstenberg-semi-transitivity. Then we provide sufficient conditions for cosine operator functions generated by adjoints of weighted composition operators to be Furstenberg-semi-transitive on the space of Radon measures on a locally compact, non-compact Hausdorff space. Also, we characterize disjoint Furstenberg-semi-transitive generalized bilateral weighted shifts on the standard Hilbert module over the C*-algebra of compact operators on a separable Hilbert space. As a special case, we consider the case when the corresponding Furstenberg family under consideration is the family of all infinite subsets of the set of natural numbers. Finally, we illustrate our results with concrete examples.
@@ 1. Introduction
The research on the dynamics of weighted composition operators has started from pioneering papers by Salas on the dynamics on weighted shifts, see [33, 35]. Since then, numerous papers on the dynamics of weighted composition operators on various function spaces have been published. In particular, the dynamics of cosine operator functions generated by weighted composition operators has been studied in for instance [10, 11, 12, 13, 14, 23, 37]. Moreover, the dynamics of cosine operator functions generated by some other operators has been investigated in for instance [22, 27, 28]. Further, Furstenberg topological transitivity of weighted composition operators, as well as of some other classes of operators, has been studied in several papers, see [2, 7, 18, 19, 29]. Finally, quite recently, topological transitivity of the adjoints of weighted composition operators has been investigated and studied in [22]. These adjoints considered in [22] act on the space of Radon measures on a locally compact, non-compact Hausdorff space. All these facts motivated us to study in this paper the Furstenberg-dynamics of cosine operator functions generated by the adjoints of weighted composition operators. In Section 3 in this paper, we introduce first the concept of $\mathcal{F}$-semi-transitivity, as a generalization of the concept of topological transitivity for supercyclicity given in [30] and as an extension of the notion of $\mathcal{F}$--transitivity Then, in Theorem 3.1, we obtain sufficient conditions for cosine operator functions generated by the adjoints of weighted composition operators to be $\mathcal{F}$-semi-transitive on the space of Radon measures (on a locally compact, non-compact Hausdorff space) equipped with the total variation norm. A special case of this theory is the case when the corresponding Furstenberg family under consideration is the family of all infinite subsets of $\mathbb{N},$ as given in Corollary 3.2. We provide also concrete examples where the sufficient conditions of Theorem 3.1 and Corollary 3.2 in this paper are satisfied, but where the sufficient conditions of [22, Proposition 3.3] are not satisfied, illustrating thus the difference between (ordinary) topological transitivity and $\mathcal{F}$-semi-transitivity of cosine operator functions.
@@ 2. Preliminaries
    If AA belongs to $\mathcal{F}$ and $B$ contains AA, then $B$ also belongs to $\mathcal{F}$.
@@ 3. $\mathcal{F}$-semi-transitive cosine operator functions
Assume that $S$ is a set and $\mathcal{F}$ is a proper Furstenberg family of subsets of $S.$
Let $(\alpha_{z})_{z\in\in\mathcal{S}}$ be a family of homeomorphisms of $\Omega$ and $(w_{z})_{z\in\in\mathcal{S}}$ be a family of positive bounded continuous functions on $\Omega$ such that $\frac{1}{w_{z}}$ is also bounded for all $z\in\mathcal{S}$. For each $z\in\mathcal{S}$, we let $T_{z}^{\ast}:=T_{\alpha_{z},w_{z}}^{\ast}$ and let $C_{z}^{\ast}:=\frac{1}{2}(T_{z}^{\ast}+S_{z}^{\ast})$. We will say that $\{C_{z}^{\ast}\}_{z\in\mathcal{S}}$ is $\mathcal{F}$-semi-transitive if for every non-empty open subsets $U_{1}$ and $U_{2}$ of $M(\Omega)$, the set
@@ 4. Disjoint $\mathcal{F}$-semi-transitive generalized weighted shifts
We start with the following definition.
Definition 4.1.
Let $S$ be a set and let $\mathcal{F}$ be a proper Furstenberg family of subsets of $S$ and $\{T_{t,1}\}_{t\in S},\dots,\{T_{t,N}\}_{t\in S}$ be families of bounded linear operators on a Banach space $X$. We say that $\{T_{t,1}\}_{t\in S},\dots,\{T_{t,N}\}_{t\in S}$ are disjoint $\mathcal{F}$-semi-transitive, or shortly $d\mathcal{F}$-semi-transitive, if for every collection of non-empty open subsets $\mathcal{O},V_{1},\dots,V_{N}$ of $X$, the set
$\{t\in S\ |\lambda_{t}\in\mathbb{R}^{+}\text{ suth that}\mathcal{O}\cap\lambda_{t}T_{t,1}^{-1}(V_{1})\cap\dots\cap\lambda_{t}T_{t,N}^{-1}(V_{N})\neq\varnothing\}$
belongs to $\mathcal{F}$.